Solve using conversion factors

A regular furnace costs $900. A high-efficiency furnace costs $1500. The average monthly bill with the
regular furnace is expected to be $80. The average monthly bill with the high-efficiency furnace is expected
to be $50. How long until the high-efficiency furnace pays for itself? Assume an interest rate of 1% per
month.
Answer #1

Data:-

a regular furnace costs $900

a high efficiency furnace costs $1500

the average monthly bill of regular is expected to be $80

the average monthly bill of high efficiency furnace is $50

interest rate of 1% per month

Solution:-

the average annual expenditure of high efficiency furnace is = (50 * 12 * 1.01) $ = 606 $

the cost of high efficiency furnace is 1500$

the total years that the high efficiency furnace pays of itself is = 1500/606 = 2.47 years